[Hallc_running] [New Logentry] Hall C RC Update (Sept 16)

brads at jlab.org brads at jlab.org
Wed Sep 16 19:00:13 EDT 2020

Logentry Text:
<p><strong>RC Daily Meetings are M-F at 4:15 PM</strong></p><pre>To join the Meeting:  <a class="external free" href="https://bluejeans.com/861439920" rel="nofollow">https://bluejeans.com/861439920</a>
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))
<br><a class="new" title="RC Update and Run plan for Sep 16,2020" href="https://hallcweb.jlab.org/wiki/index.php?title=RC_Update_and_Run_plan_for_Sep_16,2020">RC Update and Run plan for Sep 16,2020</a></pre><p><strong> Wednesday 16 Sept 2020. </strong></p><h1><span id="General_updates" class="mw-headline">General updates</span></h1><p>(See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020">d2n Run Plan 2020</a> for broader status/overview.)</p><ul><li>Lost SL 2K coldbox at 8am Tuesday morning <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848721" rel="nofollow">[1]</a>; minimum 4 hour down, cryo identified the problem but machine wasn't back up until late swing<ul><li>QWP target optimization done this down <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848806" rel="nofollow">[2]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849677" rel="nofollow">[3]</a></li!
 ><li>NMR @ ~41--2% with calibration constant 5.04 (Note: calibration constant is estimated).<ul><li>Seems to be an overall improvement of a few percent. QWP work, or due to Target Oven temperature increased from 205 to 208 at 2338 on Mon <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848466" rel="nofollow">[4]</a>?</li></ul></li></ul></li><li>Continued production running in Kin-A, Kin-X at transverse/270 (w/ T1 and T2 triggers) after standard beam profile checks during OWL</li><li>Moller measurement planned for Wednesday DAY starting at 8:30am<ul><li>Hall A's Moller will ran Wednesday DAY as well</li><li>Hall C result: 85.4% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849516" rel="nofollow">[5]</a>; (Hall A: ~87% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849466" rel="nofollow">[6]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849477" rel="nofollow">[7]</a>)</li></ul></l!
 i><li>HMS dipole tripped at 8:10am <a class="external autonumb!
 er" href="https://logbooks.jlab.org/entry/3849359" rel="nofollow">[8]</a>(Current lead mass flow fault triggered by unstable recovery pressure)<ul><li>HMS dipole PID loop adjusted by Heidi <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849434" rel="nofollow">[9]</a></li></ul></li><li>Setup spectrometers to Kin-C, Kin-Z (Target stays at transverse/270) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849410" rel="nofollow">[10]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849418" rel="nofollow">[11]</a></li><li>Production running in Kin-C, Kin-Z at transverse/270 (T1 trigger only) resumed around 1pm</li><li>We'll continue production data taking here until switch to 1-pass at 6am on Friday morning.</li></ul><p> </p><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2><ul><li><a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2!
 020#1-pass_Running">1-pass running</a>transition scheduled for 6am Friday, Sept 18<ul><li>RadCon is aware of the plans</li><li>Hall A will setup the charge-asym feedback system for us (P. King, et al)\</li><li>Junhao has some estimates of the NMR calibration coeffs for 1-pass 0 and 180 settings <a class="external autonumber" href="https://logbooks.jlab.org/entry/3849028" rel="nofollow">[12]</a></li></ul></li></ul><ul><li>Melanie, Murchhana, Chao, Sylvester: please check good electron yields per Coulomb for Kin-A/X runs with T1 vs. T3 (T2)<ul><li>Mostly complete. Early results don't show any surprises..</li></ul></li></ul><ul><li>EPR calibrations cross checks [Melanie]<ul><li>So far it all looks good, checks ongoing</li><li>Summary log entries: EPR(270, Kin-Z) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3847972" rel="nofollow">[13]</a>, EPR(90, Kin-X) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848215" rel="nofollow">[14]</a!
 ></li></ul></li></ul><ul><li>Bill, Murchhana check compass/field measur!
 ements<ul><li>Also checking with Jixie re: modeled fields</li><li>Need to make a write-up that clearly defines all the different coordinate systems used, swapped power supply/polarity issues, etc.</li></ul></li></ul><ul><li>Start prepping for post-run work<ul><li>See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#POST-run_Calibrations.2FMeasurements">POST-run Calibrations/Measurements</a> for list</li></ul></li></ul><h2><span id="Pending_Issues" class="mw-headline">Pending Issues</span></h2><ul><li>Review and do any final prep/planning for 1-pass running<ul><li>Paddle optimization for 3He Elastic xsec running [Chao?]</li><li>EPR calibration for 1-pass 8.5deg setting (estimate from prior calibration?)</li></ul></li></ul><ul><li>Beam energy drifts<ul><li>Should confirm that MCC remembers to shift the energy lock to Hall C when Hall A turns off</li><li>Just something to keep an eye on. Has not been a problem lately.</li><li>Hall C!
  5-pass polarization goes down as energy (HALLC:p) goes up <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844493" rel="nofollow">[15]</a></li></ul></li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3849805
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200916/37a06fc0/attachment.html>

More information about the Hallc_running mailing list